Operations Report – December 10th

Crew 184 Operations Report 10/12/2017

SOL: 9

Name of person filing report: HuntJ

Non-nominal systems: Generator non-functional, repaired and returned to operations in afternoon. (Operational Again!!!)

Notes on non-nominal systems: Mission Support found a solution with the generator today. It turns out that the ground and positive connectors to the battery were loose. Tightening them and a voltimeter check proved the battery was just fine. One more attempt at using jumper cables and the generator is back and running! Forward plan now is to leave it on until tomorrow morning (with fingers crossed that it stays on through the night). A decision will be made tomorrow morning on whether to turn the generator back off or leave it running.

Generator (hours run): Back on around 2:00pm. Running 3 hours and still going!

Solar— SOC @ 4:00pm: 99%

Diesel – 100%

Propane – 70 Percent Volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1.5 Gallons

Water (trailer) – 600. gallons

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 47 gallons

Toilet tank emptied: No

ATV’s Used: None,

Oil Added: No

ATV Fuel Used: None

# Hours the ATVs were Used today: 00:00 Hours

Notes on ATVs: No ATVs used today but were turned on for status checks. All nominal except for the Yamaha 300 which was determined a few days ago to have a dead battery.

Deimos rover used: Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 88.8%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 01:00

Beginning charge: 100%

Ending charge: 51%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 23%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 00:30 Hours

Beginning charge: 100%

Ending charge: 59%

Currently charging: Yes

HabCar used and why, where? NO

General notes and comments: Power is back and we’re back in sim!!

Summary of internet: 1.0 GB of Bandwidth

Summary of suits and radios: Radios are nominal after getting charged again.

EVA Suits:

EVA 6: Fan does not turn on.

EVA 8: Latches that secure the back panel are loose. Engaging them again does not change this. Concern is that back panel would fall off during EVA.

We will try to troubleshoot some of these issues and will update with changes.

Summary of Hab operations: Nominal

Summary of GreenHab operations: See Greenhab report

Summary of ScienceDome operations: Currently being cleaned

Summary of RAM operations: Not Operational

Summary of health and safety issues: All Healthy

Questions, concerns and requests to Mission Support:

Request more ATV Fuel. Can probably get through 2 more EVAs before exhausting remaining reserves.
Non burnable trash in airlock ready for pickup.

Copyright © The Mars Society. All rights reserved. | Main Site